OpenStudy (anonymous):

d/dx(integral cos(t^2)dt)

OpenStudy (frostbite):

\[\frac{ d }{ dx }\int\limits_{}^{}\cos(t ^{2})dt\] right?

OpenStudy (frostbite):

If that is the thing you differentiate due to a constant and then the whole thing become 0.

OpenStudy (loser66):

I don't think so, it is= cos (t^2)

OpenStudy (anonymous):

yes it is ddx∫cos(t2)dt

OpenStudy (anonymous):

how do you find the definite integral

OpenStudy (loser66):

by definition, derivative cancel out with integral and you get the original function without calculating any thing

OpenStudy (anonymous):

oh alright. Thank you.

OpenStudy (loser66):

need more explanation? need more example?

OpenStudy (anonymous):

so basically you are saying that d/dx and integral cancel which leaves you with the original equation of cos (t^2) dt right?

OpenStudy (loser66):

no dt. just cos (t^2)

OpenStudy (loser66):

for example: if